# Pólya’s conjecture up to $\epsilon$-loss and quantitative estimates for the remainder of Weyl’s law 00footnotetext: 2020 Mathematics Subject Classification. Primary 35P15; Secondary 35P20, 42B37. Key words and phrases: Dirichlet eigenvalue, Weyl law, Pólya conjecture

### v1

@@ 3 Pólya’s conjecture on a class of domains
Let us start with proving the qualitative result regarding removing general piecewise smooth set.
@@
Let $\Omega=\Omega_{1}\times\Omega_{2}\setminus\overline{\Omega_{3}}$. Up to a scaling, we may assume that $|\Omega_{3}|=1$. It follows from Seeley [32, 33] (see also [18, Corollary 17.5.11 & (17.5.16)’]) that for some $\Lambda>0$, it holds for $\lambda>\Lambda$ that
@@ 3.1 The case of two dimension
Note that for a given $\lambda_{k}$, $\mathcal{N}_{Q}^{D}(\lambda_{k})$ is dominated by the number of unit cubes contained in the first quadrant intersecting the ball $B(0,\sqrt{\lambda_{k}/\pi})$.
So it is readily to see that for all $k\in{\mathbb{N}}$ that $\lambda_{k}\geq 4\pi k.$ In fact Theorem 2.3 gives a better result. We however shall need the opposite estimates.
@@
Note that for $\lambda\leq 2\pi^{2}/\ell(Q)^{2}$,
$\frac{\ell(Q)^{2}}{4\pi}\lambda-\frac{\sqrt{2}\ell(Q)}{2}\sqrt{\lambda}<0<\mathcal{N}_{Q}^{D}(\lambda),$
@@
Moreover, by Lemma 3.2 [(ii)], we see that for $\lambda\geq\lambda_{k_{0}}(\Omega)$,
$\mathcal{N}^{D}_{\Omega\cup\cup_{k\in{\mathbb{N}}}Q_{k}}(\lambda)$ $=\mathcal{N}^{D}_{\Omega}(\lambda)+\sum_{k\in{\mathbb{N}}}\mathcal{N}_{Q_{k}}(\lambda)$
@@ 3.2 Three and higher dimensions
We next provide the proof for $n\geq 3$. Let $Q\subset{\mathbb{R}}^{n}$ be a unit cube. Similar to Lemma (3.2), we have
@@
since $\sqrt{\lambda}>\pi\sqrt{3}/\ell(Q)$. For $\sqrt{\lambda}\leq\pi\sqrt{3}/\ell(Q)$, note that
@@
Since $|\Omega_{1}|\geq 9\sqrt{3}\mathscr{S}_{Q}\geq 9\sqrt{3}\mathscr{V}_{Q}$ and $|\Omega_{2}|>1$, it holds that
@@
Since $\lambda_{k}(\Omega)>61$, $|\Omega_{2}|>1$, we have that
@@
$\mathcal{N}^{D}_{\Omega\cup\Omega_{3}}(\lambda)$ $=\mathcal{N}^{D}_{\Omega}(\lambda)+\sum_{j\in{\mathbb{N}}}\mathcal{N}_{Q_{j}}(\lambda)$
@@
$\mathcal{N}^{D}_{\Omega\cup\Omega_{3}}(\lambda)$ $=\mathcal{N}^{D}_{\Omega}(\lambda)+\sum_{j\in{\mathbb{N}}}\mathcal{N}_{Q_{j}}(\lambda)$
@@ 3.3 Some other class satisfying Pólya’s conjecture
The required inequalities followed from the Rayleigh-Faber-Krahn inequality ($\lambda_{1}(\Omega)$) and the Krahn-Szegö inequality ($\lambda_{2}(\Omega)$), see [16].
Note that $\Omega\cup\cup_{j}Q_{j}\subsetneq\Omega_{1}$, and $|\Omega\cup\cup_{j}Q_{j}|=|\Omega_{1}|$. By Lemma 3.5 and Lemma 3.5 (ii), and the fact $\Omega$, $Q_{j}$ are mutually disjoint, we find that
$\mathcal{N}^{D}_{\Omega_{1}}(\lambda)$ $\geq\mathcal{N}^{D}_{\Omega\cup\cup_{j}Q_{j}}(\lambda)=\mathcal{N}^{D}_{\Omega}(\lambda)+\sum_{j\in{\mathbb{N}}}\mathcal{N}_{Q_{j}}(\lambda)$
@@
Inspired by [15, Theorem 1.2], using Theorem 3.7 we have another class of domains satisfying Pólya’s conjecture.
@@
4 Proof of the main results
In this section, we prove our main uniform estimates on Lipschitz domains. Let us recall the Whitney decomposition, see [14, Appendix J].
@@ 4 Proof of the main results
Step 2. Lower bound for the Dirichlet eigenvalues. To prove the lower bound for the Dirichlet eigenvalues of $\Omega$, let us set $\Omega_{1}=\mathscr{R}\setminus\overline{\Omega}$. Obviously, $\Omega_{1}$ is also a Lipschitz domain. Moreover, it holds that
@@
Remark 4.2.
@@
$\quad+\frac{\omega(n)}{(2\pi)^{n}}\lambda_{k}(\Omega)^{\frac{n-1}{2}}\left(|\partial\mathscr{R}\setminus\partial\Omega|\left(\frac{\pi r_{in}(\Omega_{1})}{|I_{min}|}+5n^{2}\pi\log_{2}\frac{10|I_{min}|\sqrt{\lambda_{k}(\Omega)}}{\pi}\right)-C_{1}(n-1)|\mathscr{R}_{n-1}|\right).$
Obviously, |Ω||($|\partial\mathscr{R}\setminus\partial\Omega|\leq|\partial(\mathscr{R}\setminus\overline{\Omega})|$. The number $\Lambda(\epsilon,\Omega)$ from Theorem 1.3 can be taken smaller than (1.7) as
$\frac{2n\mathrm{diam}(\Omega)^{n-1}}{|\Omega|\sqrt{\Lambda(\epsilon,\Omega)}}\left(\frac{\pi}{2}-\frac{C_{1}(n-1)}{2n}+5n^{2}\pi\log_{2}\frac{10\mathrm{width}(\Omega)\sqrt{\Lambda(\epsilon,\Omega)}}{\pi}\right)=\epsilon,$
where $C_{1}(n-1)$ is as in (1.9), see the following proof of Theorem 1.3.
@@
$\quad+\frac{\omega(n)}{(2\pi)^{n}}\lambda_{k}(\Omega)^{\frac{n-1}{2}}\left(C_{Lip}(\Omega)|\partial(\mathscr{R}\setminus\Omega)|\left(\frac{\pi r_{in}(\mathscr{R}\setminus\Omega)}{|I_{min}|}+5n^{2}\pi\log_{2}\frac{10|I_{min}|\sqrt{\lambda_{k}(\Omega)}}{\pi}\right)-C_{1}(n-1)|\mathscr{R}_{n-1}|\right)$
@@
The proof is complete. ∎
For open sets, we also have a uniform estimate. Let us remark that, since the boundary of an open set may be complicate, in general one can not have a uniform estimate in term of its boundary as in Theorem 1.5. Note that $C_{1}(n-1)$ is defined via (1.9).
@@ Acknowledgments
R. Jiang wishes to thank Prof. Xiao Zhong for explaining to him Pólya’s conjecture and helpful discussions at University of Jyväskylä around 2010. R. Jiang was partially supported by NNSF of China (12471094 & 11922114), F.H. Lin was in part supported by the National Science Foundation Grant DMS2247773 and DMS1955249.

### v2

@@ 3 Pólya’s conjecture on a class of domains
Let us start with proving the qualitative result regarding removing general piecewise smooth set. Theorem 1.7 follows immediately from the following result.
@@
Let $\Omega=\Omega_{1}\times\Omega_{2}\setminus\overline{\Omega_{3}}$. Up to a scaling, we may assume that $|\Omega_{3}|=1$. It follows from Seeley [35, 36] (see also [20, Corollary 17.5.11 & (17.5.16)’]) that for some $\Lambda>0$, it holds for $\lambda>\Lambda$ that
@@
Remark 3.2.
Note that (3.1) can be stated for all $\lambda>0$ without referring to $\Lambda$, but generally with larger constant $C(\Omega_{3})$. We state in this way to show that the combined argument, that is for small eigenvalues using Corollary 2.4 while for large constant using the subtracting argument, will give better estimate, see Remark 3.5 below.
@@ 3.1 The case of two dimension
Note that for a given $\lambda_{k}$, $\mathcal{N}_{Q}^{D}(\lambda_{k})$ is dominated by the number of unit cubes contained in the first quadrant intersecting the ball $B(0,\sqrt{\lambda_{k}}/\pi)$. So it is readily to see that for all $k\in{\mathbb{N}}$ that $\lambda_{k}\geq 4\pi k.$ In fact Theorem 2.3 gives a better result. We however shall need the opposite estimates.
@@
Noting that for $\lambda\leq 2\pi^{2}/\ell(Q)^{2}$,
$\frac{\ell(Q)^{2}}{4\pi}\lambda-\frac{\sqrt{2}\ell(Q)}{2}\sqrt{\lambda}\leq 0\leq\mathcal{N}_{Q}^{D}(\lambda),$
@@
Moreover, by Lemma 3.3 (ii), we see that for $\lambda\geq\lambda_{k_{0}}(\Omega)$,
$\mathcal{N}^{D}_{\Omega\cup\cup_{k\in{\mathbb{N}}}Q_{k}}(\lambda)$ $=\mathcal{N}^{D}_{\Omega}(\lambda)+\sum_{k\in{\mathbb{N}}}\mathcal{N}^{D}_{Q_{k}}(\lambda)$
@@ 3.2 Three and higher dimensions
We next provide the proof for $n\geq 3$. Similar to Lemma (3.3), we have
@@
for $\sqrt{\lambda}>\pi\sqrt{3}/\ell(Q)$. For $\sqrt{\lambda}\leq\pi\sqrt{3}/\ell(Q)$, note that
@@
Since $|\Omega_{1}|\geq 9\sqrt{3}\mathscr{S}_{Q}\geq 9\sqrt{3}\mathscr{V}_{Q}$ and $|\Omega_{2}|\geq 1$, it holds that
@@
Since $\lambda_{k}(\Omega)>61$, $|\Omega_{2}|\geq 1$, we have that
@@
$\mathcal{N}^{D}_{\Omega\cup\Omega_{3}}(\lambda)$ $=\mathcal{N}^{D}_{\Omega}(\lambda)+\sum_{j\in{\mathbb{N}}}\mathcal{N}^{D}_{Q_{j}}(\lambda)$
@@
$\mathcal{N}^{D}_{\Omega\cup\Omega_{3}}(\lambda)$ $=\mathcal{N}^{D}_{\Omega}(\lambda)+\sum_{j\in{\mathbb{N}}}\mathcal{N}^{D}_{Q_{j}}(\lambda)$
@@ 3.3 Some other class satisfying Pólya’s conjecture
The required inequalities for $k\leq 2$ followed from the Rayleigh-Faber-Krahn inequality ($\lambda_{1}(\Omega)$) and the Krahn-Szegö inequality ($\lambda_{2}(\Omega)$), see [18].
Let us consider $k\geq 3$. Note that $\Omega\cup\cup_{j}Q_{j}\subsetneq\Omega_{1}$, and $|\Omega\cup\cup_{j}Q_{j}|=|\Omega_{1}|$. By Lemma 3.6 and Lemma 3.6 (ii), and the fact $\Omega$, $Q_{j}$ are mutually disjoint, we find that
$\mathcal{N}^{D}_{\Omega_{1}}(\lambda)$ $\geq\mathcal{N}^{D}_{\Omega\cup\cup_{j}Q_{j}}(\lambda)=\mathcal{N}^{D}_{\Omega}(\lambda)+\sum_{j\in{\mathbb{N}}}\mathcal{N}^{D}_{Q_{j}}(\lambda)$
@@
Inspired by [17, Theorem 1.2], using Theorem 3.8 we have another class of domains satisfying Pólya’s conjecture.
@@
4 Pólya’s conjecture up to $\epsilon$-loss and quantitative remainder estimate
In this section, we prove our main uniform estimates on Lipschitz domains. Let us recall the Whitney decomposition, see [16, Appendix J].
@@ 4 Pólya’s conjecture up to $\epsilon$-loss and quantitative remainder estimate
Step 2. Lower bound for the Dirichlet eigenvalues. To prove the lower bound for the Dirichlet eigenvalues of $\Omega$, let us set $\Omega_{1}=\mathscr{R}\setminus\overline{\Omega}$. Obviously, $\Omega_{1}$ is also a Lipschitz domain. Moreover, by (1.6), it holds that
@@
We next sketch the proof of Corollary 1.6.
Proof of Corollary 1.6.
@@
$\quad+\frac{\omega(n)}{(2\pi)^{n}}\lambda_{k}(\Omega)^{\frac{n-1}{2}}\left(|\partial\mathscr{R}\setminus\partial\Omega|\left(\frac{\pi r_{in}(\mathscr{R}\setminus\overline{\Omega})}{|I_{min}|}+5n^{2}\pi\log_{2}\frac{10|I_{min}|\sqrt{\lambda_{k}(\Omega)}}{\pi}\right)-C_{1}(n-1)|\mathscr{R}_{n-1}|\right).$
The proof is complete. ∎
@@
$\quad+\frac{\omega(n)}{(2\pi)^{n}}\lambda_{k}(\Omega)^{\frac{n-1}{2}}\left(C_{Lip}(\Omega)|\partial(\mathscr{R}\setminus\overline{\Omega})|\left(\frac{\pi r_{in}(\mathscr{R}\setminus\overline{\Omega})}{|I_{min}|}+5n^{2}\pi\log_{2}\frac{10|I_{min}|\sqrt{\lambda_{k}(\Omega)}}{\pi}\right)-C_{1}(n-1)|\mathscr{R}_{n-1}|\right)$
@@
If $\Omega$ is convex, then applying Corollary 1.6 gives that
$k$ $\leq\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda_{k}(\Omega)^{n/2}\left(1+\frac{2n\mathrm{diam}(\Omega)^{n-1}}{\sqrt{\lambda_{k}}|\Omega|}\left(\frac{\pi}{2}-\frac{C_{1}(n-1)}{2n}+5n^{2}\pi\log_{2}\frac{10\mathrm{width}(\Omega)\sqrt{\lambda_{k}}}{\pi}\right)\right).$
The remaining proof is the same. The proof is complete. ∎
For open sets, we also have a uniform estimate. Let us remark that, since the boundary of an open set may be complicate, in general one can not have a uniform estimate in term of its boundary as in Theorem 1.5; see [9, Theorem 3.14] for similar but earlier result.
Note that $C_{1}(n-1)$ is defined via (1.11).
@@ Acknowledgments
We wish to thank Prof. Rupert Frank and Prof. Simon Larson for bringing references [9, 12, 28] to us and several useful comments. R. Jiang wishes to thank Prof. Xiao Zhong for explaining to him Pólya’s conjecture and helpful discussions at University of Jyväskylä around 2010. R. Jiang was partially supported by NNSF of China (12471094 & 11922114), F.H. Lin was in part supported by the National Science Foundation Grant DMS2247773 and DMS1955249.